A 10.0−mL solution of 0.330 M NH3 is titrated with a 0.110 M HCl solution. Calculate the pH after the following additions of the HCl solution:

(a) 0.00 mL



(b) 10.0 mL



(c) 30.0 mL



(d) 40.0 mL
Answer #1

Given that The volume ofNH3= 10.0 ml The molarity of NH3 0.3 M The number of moles of NH3 Molarity * volume in liters =0.314

a) The volume of HCI 0.0 ml There fore in this case the solution contains only ammonia We know that Where C-concentration of

b) The volume of HC 10.0 ml The number of moles of HCl - molarity * volume of HCl in liters -0.1 M 0.01 L 0.001 mo1 The numbe

According to the Henderson - Hasselbalch equation salt [base] INH3] [0.002mol] [0.001mol] 4.74+log 4.74-0.301 - 4.4389

pH-14-рон 14-4.4389 9.561 1

c) The volume of HCl= 20.0 ml The number of moles of HCl = molarity * volume of HCl in liters -0.1 M 0.02L 0.002 mol The numb

According to the Henderson - Hasselbalch equation pOH = pKb +log [salt] [acid] =pKb +log [NH3] = 4 74 + logt0001.nol] 4.74+0.

pH- 14- pOH = 14-5.041 = 8.959

d) The volume of HCI 30.0 ml It is exactly equivalence point so at this point there is only NH4 ion and Cr ion are in solutio

Ka of NH4 ion-Kw /Kb of NH3 1x1014 1.8x10-3 556x10-ю We know that 15.56x10-10 x 0.075M 6.4576x 106 M pH- log(H -log (6.4576x

e) The volume of HCl 40.0 ml In this case the solution contains excess of HC1 so we have to calculate the concentration of ex